# 11footnotetext: Implementation of AQUA is public available at https://github.com/tychenn/AQUA

### v1

@@ Abstract
As Retrieval-Augmented Generation (RAG) evolves into service-oriented platforms (Rag-as-a-Service) with shared knowledge bases, protecting the copyright of contributed data becomes essential. Existing watermarking methods in RAG focus solely on textual knowledge, leaving image knowledge unprotected. In this work, we propose AQUA, the first watermark framework for image knowledge protection in Multimodal RAG systems. AQUA embeds semantic signals into synthetic images using two complementary methods: acronym-based triggers and spatial relationship cues. These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, being efficient, effective and imperceptible. Experiments across diverse models and datasets show that AQUA enables robust, stealthy, and reliable copyright tracing, filling a key gap in multimodal RAG protection.
@@ 1 Introduction
Figure 1: Overview of the RAG-as-a-Service (RaaS) workflow. Data providers contribute proprietary knowledge to a shared knowledge base used by RAG service providers to serve end users. Data providers can issue watermark probe queries to RAG services. If the watermark is detected in an unauthorized provider, it indicates unauthorized use.
Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks, but they often suffer from hallucinations and outdated knowledge learned in the static parameters. To mitigate these limitations, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023a) has emerged as a promising paradigm that augments LLMs with up-to-date external knowledge retrieved at inference time. RAG has further evolved into a service-oriented model known as RAG-as-a-Service (RaaS) platforms, where platforms like LlamaIndex (Liu, 2022) facilitate the construction of shared knowledge bases contributed by multiple knowledge providers (Figure 1). Importantly, these systems adopt a “usable but not visible” policy: RAG service provider can leverage the contributed knowledge without directly accessing the raw data.
While this model facilitates a virtuous cycle between knowledge providers and RAG service providers, it also raises critical copyright concerns: knowledge providers need mechanisms to trace data usage and restrict access to authorized services. Since unauthorized RAG providers typically utilize the entire shared knowledge base, knowledge provider can embed watermarks at the knowledge base level. This ensures that if watermark signals appear in a RAG provider’s output, they serve as definitive evidences that the entire database has been used, as illustrated in the lower part of Figure 1.
@@
Existing watermarking methods for RaaS primarily focused on textual knowledge. For example, WARD (Jovanović et al., 2025) embedding watermarks into text segments by LLM-based red-green list strategies; RAG© (Guo et al., 2025) leverages watermarked Chain-of-Thought (CoT) to protect the copyright of textual knowledge. However, these methods are modality-specific, limited to text modality and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities. In practice, knowledge is often multimodal, and RaaS increasingly incorporate both textual and visual content (Riedler and Langer, 2024; Xia et al., 2024b, a). This exposes a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS. To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where generator integrates retrieved image knowledge and textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022a; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
@@
Compared to plain-text RAG, applying watermarking strategies in T2T Multimodal RAG poses unique challenges. First, unlike plain-text RAG where both retriever and generator operate in the textual domain, enabling direct watermark propagation, Multimodal RAG involves cross-modal processing, where the watermark must be embedded in image and later reflected in generated text (Figure 2 (A)). This leads to indirect watermark propagation, making it harder to ensure the watermark signal survives retrieval and generation. Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), which makes watermark images harder to be consistently retrieved by probe queries compared to plain-text RAG. Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) typically add imperceptible perturbations onto the image using optimization-based methods, which are implicit and designed for detection directly on the image itself. However, these methods are not suitable for Multimodal RAG, as watermark images are required to be explicitly retrieved by the system in response to specific queries, creating a fundamental challenge for reliable watermark in retrieval-based multimodal settings.
@@
To address above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG. Specifically, AQUA watermarking framework includes two complementary watermarking methods: $\emph{AQUA}_{acronym}$ and $\emph{AQUA}_{spatial}$. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images. In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of generators (Vision-Language Models (VLMs) (Achiam et al., 2023; Team et al., 2023; Huang et al., 2023) ) to generate detectable textual response: the full name of the acronyms. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. For models with limited OCR ability, $\emph{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g. unusual positional relationships), and leverage generators’ understanding of spatial semantics to answer position-related probe queries. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator. Both methods introduces semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate. Together, these two methods provide a flexible, robust solution to the unique challenges of watermarking in Multimodal RAG systems, supporting both black-box and white-box deployments.
@@
Despite simplicity, our novel insights of using synthetic images with special acronyms texts and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG. We evaluate AQUA across diverse Multimodal RAG and datasets spanning different domains. The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevent false positives from common image content, (3) remain imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to common image transformations.
Our contribution can be summarized as follow:
@@
  • We propose AQUA, the first watermarking framework tailored for image knowledge copyright protection in Multimodal RAG systems, addressing indirect watermark propagation, and successful retrieval under unapparent distribution shifts and explicit watermark injection;
@@
  • We design two complementary watermarking strategies, $\emph{AQUA}_{acronym}$, $\emph{AQUA}_{spatial}$ to support more realistic black-box scenarios;
  • We perform comprehensive experiments on two well-known datasets, utilizing four prevalent pretrained VLMs (LLaVA-NeXT, InternVL3, Qwen-VL-Chat and Qwen2.5-VL-Instruct) to assess the effectiveness, harmlessness, stealthiness and robustness of AQUA.
@@ 2.1 Multimodal Retrieval-Augmented Generation
Plain-text Retrieval-Augmented Generation (RAG). (Lewis et al., 2020; Singh et al., 2021; Wang et al., 2023; Asai et al., 2023b; Xu et al., 2024; Zhao et al., 2024; Tan et al., 2025) proposed several RAG frameworks to address the hallucination in LLMs. They mainly focus on textual external knowledge, and these methods are inadequate to process and integrate non-textual modalities inherent in real-world information.
Multimodal Retrieval-Augmented Generation (Multimodal RAG). (Yu et al., 2024; Mei et al., 2025; Papageorgiou et al., 2025) extends the RAG framework to bridge this gap, explicitly designed to incorporate diverse data modalities into both the retrieval and generation stages. A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g. CLIP (Radford et al., 2021) ), to map different modalities (e.g., text and images) into a shared semantic embedding space. This unification allows standard vector search algorithms like cosine similarity to retrieve relevant items across modalities based on semantic relatedness.
@@ 2.2 RAG Watermarking
Several watermarking approaches have been proposed to protect the copyright of textual knowledge in RAG. WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all the texts in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024) . RAG-WM (Lv et al., 2025) presents a black-box RAG watermarking approach that leverages interactions among multiple LLMs to generate high-quality watermarks. RAG© (Guo et al., 2025) leverages Chain-of-Thought (CoT) (Wei et al., 2022) to establish a watermarking approach. DMI-RAG (Liu et al., 2025) performs dataset membership inference by injecting a small number of synthetic, watermarked "canary" documents into the Intellectual Property (IP) dataset. However, existing methods on watermarking knowledge base in RAG system have exclusively focused on purely textual data. To the best of our knowledge, no prior work has addressed the protection of knowledge copyright in Multimodal RAG systems, particularly those integrating image and text modalities, via watermarking techniques.
@@ 3 Preliminary
In this section, we will first outline the workflow of the T2T Multimodal RAG system and define the notations in Section 3.1. Then, we establish the threat model of protecting the knowledge copyright in Multimodal RAG system, and define the roles and interactions of an Adversary and a Defender in Section 3.2.
@@ 3.1 Multimodal RAG System Workflow
The T2T Multimodal RAG system contains three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$. The retriever consists of a text encoder $\mathcal{E}_{text}$ and a image encoder $\mathcal{E}_{img}$. Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are pre-processed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
Knowledge Retrieval. The retriever accepts the user’s text query $T$ as input, and process it into the same latent space as image: $e_{T}$: $e_{T}=\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$. Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):=\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$. Based on these similarity scores $s_{i}$, the retriever selects the top-k most relevant images as output:
@@
Augmented Generation. The original text query $T$ and the retrieved set of images $D_{retrieved}$ are combined and passed to the generator $\mathcal{G}$ to produce the final answer: $A=\mathcal{G}(D_{retrieved},T)$
@@ 3.2 Threat Model
We consider the image knowledge copyright protection in Multimodal RAG service.
Defender represents the knowledge provider, aiming to detect and prevent unauthorized use of their proprietary image knowledge by external Multimodal RAG services. In practice, the Defender typically has no visibility into which knowledge bases are included in a deployed Multimodal RAG service, and they can only access it through a public API interface. Defender can only operate on their own datasets to implement protection mechanisms such as injecting watermarks before contributing their data to a RaaS.
Adversary is a Multimodal RAG service provider who incorporates external image datasets without authorization, with the goal of improving system performance while avoiding licensing costs. The Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, which creates an opportunity for Defender to audit its misuse.
@@ 4 Methodology
AQUA is a watermarking framework designed to protect the image knowledge copyrights in Multimodal RAG service, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness. In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textit{AQUA}_{acronym}$ and $\textit{AQUA}_{spatial}$. For each method, we will first introduce the principle of designing watermarks and then clarify how this method can be verified in statistical strategies.
Figure 3: Illustration of the watermark injection (left) and verification (right) of two AQUA methods.
4.1 $\textit{AQUA}_{arconym}$
Watermark Injection. $\textit{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images. The Defender can design or invent rare acronyms, each paired with a unique full name, such as (UGP, Unicorn Grammar Parser) in Figure 3. Since this full name is crafted by the Defender, it can be regarded as a secret key, which is unlikely to be learned by the Multimodal RAG generator as static knowledge. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. The acronym pair can also be generated in large quantities using LLM (e.g., Gemini-2.5-Pro), with the ability of In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are relegated in Appendix A.2. Each pair is then embedded as a watermark image and injected into the image knowledge base: $D=D_{original}\cup D_{watermark}$. These images are designed to be minimally invasive and do not affect the model’s utility for normal queries.
Watermark Verification. In verification phase, these acronyms are decoded through the OCR ability of generator, to generate detectable textual responses: the full name of the acronyms. Each watermark image has its own probe query $T_{probe}$ which can be used by the knowledge provider to detect unauthorized use. The $T_{probe}$ consists of two parts: a trigger $T_{trigger}$, used by the retriever to retrieve the watermark images, and an instruction $T_{instruction}$, which prompts the generator to generate the watermark-included responses that can be detected. We can formulate this construction as: $T_{probe}=T_{tigger}\oplus T_{instruction}$. For examples, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”. Following (Wu et al., 2024; Ha et al., 2025) to verify the watermark signal, we define a strict exact match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$:

### v2

# arXiv:2506.10030v2[cs.CR] 28 Feb 2026

@@ Abstract
As Retrieval-Augmented Generation (RAG) evolves into service-oriented platforms (RAG-as-a-Service) with shared knowledge bases, protecting the copyright of contributed data becomes essential. Existing watermarking methods in RAG focus solely on textual knowledge, leaving image knowledge unprotected. In this work, we propose AQUA, the first watermark framework for image knowledge protection in Multimodal RAG systems. AQUA embeds semantic signals into synthetic images using two complementary methods: acronym-based triggers and spatial relationship cues. These techniques ensure watermark signals survive indirect watermark propagation from image retriever to textual generator, and are efficient, effective, and imperceptible. Experiments across diverse models and datasets show that AQUA enables robust, stealthy, and reliable copyright tracing, filling a key gap in Multimodal RAG protection. The implementation of AQUA is publicly available at https://github.com/tychenn/AQUA.
@@ 1 Introduction
Figure 1: Overview of the RAG-as-a-Service (RaaS) workflow. Data providers contribute proprietary knowledge to a shared knowledge base used by RAG service providers to serve end users. Data providers can issue watermark probe queries to RAG services. If the watermark is detected in an unauthorized RAG service, it indicates unauthorized use.
Large Language Models (LLMs) often suffer from hallucinations and outdated knowledge, which Retrieval-Augmented Generation (RAG) mitigates by retrieving external knowledge at inference time (Lewis et al., 2020; Guu et al., 2020; Asai et al., 2023). RAG has further evolved into RAG-as-a-Service (RaaS), where platforms such as LlamaIndex (Liu, 2022) enable shared knowledge bases contributed by multiple providers (Figure 1). These systems follow a “usable but not visible” policy: service providers can use contributed knowledge without direct access to raw data.
While RaaS enables a mutually beneficial ecosystem between knowledge providers and service platforms, it also introduces copyright and ownership challenges. In particular, providers require mechanisms to reliably trace data usage and restrict access to authorized services. Since unauthorized RAG providers typically utilize the entire shared knowledge base, a practical solution is to embed watermarks at the knowledge base level. The detection of these watermark signals in a provider’s output can then serve as strong evidence of unauthorized data usage (Figure 1, bottom).
Existing watermarking methods for RaaS have primarily focused on textual knowledge (Jovanović et al., 2025; Guo et al., 2025). However, these methods are modality-specific, are limited to the text modality, and cannot be directly applied to non-textual knowledge due to the distinct characteristics of other modalities. In practice, RaaS systems increasingly integrate multimodal knowledge, combining textual and visual content (Riedler and Langer, 2024; Xia et al., 2024a; b). This creates a fundamental gap and leaves a critical vulnerability in the copyright protection of Multimodal RaaS. To address this gap, we focus on a representative subclass: text-to-text (T2T) Multimodal RAG, where the generator integrates retrieved image knowledge and the textual query to generate textual responses (Yasunaga et al., 2022; Chen et al., 2022; Lin and Byrne, 2022; Sun et al., 2024; Zhu et al., 2024).
@@
Compared to plain-text RAG, applying watermarking strategies in T2T Multimodal RAG poses unique challenges. First, while text-based RAG supports direct watermark propagation, Multimodal RAG requires embedding the watermark in images and reflecting it in generated text, resulting in indirect propagation that is harder to preserve. Second, unlike textual watermarks typically involving unusual tokens resulting in obvious distribution shift from original knowledge (Chen et al., 2024c; Cheng et al., 2024), image knowledge differs only at the pixel level while preserving semantic naturalness, resulting in unapparent distribution shifts (Figure 2 (B)), that reduce retrievability. Moreover, existing image watermarking methods (Luo et al., 2020; Chen et al., 2024a) rely on implicit perturbations designed for image-level detection, but Multimodal RAG requires watermarks to be explicitly retrieved through queries, making them unsuitable for retrieval-based multimodal settings.
@@
To address the above challenges in image knowledge copyright protection, we propose AQUA, a novel watermarking framework tailored for T2T Multimodal RAG. Specifically, the AQUA watermarking framework includes two complementary watermarking methods: $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$. $\textbf{AQUA}_{acronym}$ addresses indirect watermark propagation by embedding uncommon acronyms and their full names into synthetic images. In the verification phase, these acronyms are decoded through the Optical Character Recognition (OCR) abilities of Vision-Language Models (VLMs) (Achiam et al., 2023; Gemini Team et al., 2023; Huang et al., 2023) to generate detectable textual responses: the full name of the acronyms. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing.
For models with limited OCR ability, $\textbf{AQUA}_{spatial}$ is designed to create synthetic images with special object configurations (e.g., uncommon positional relationships), and to leverage generators’ understanding of spatial semantics to answer position-related probe queries. These positional relationships can bridge the gap between image semantics and textual outputs, allowing indirect watermark propagation from retriever to generator. Both methods introduce semantic distinctiveness by embedding subtle semantic cues into natural-looking images, allowing explicit retrieval while maintaining a high retrieval rate. Together, these two methods provide a flexible, robust solution to the unique challenges of watermarking in Multimodal RAG systems, supporting both black-box and white-box deployments.
@@
Despite its simplicity, our novel insights of using synthetic images with special acronyms text and special positional relationships as watermark carriers are particularly effective and efficient in bridging the gap between image-based watermarking and textually detectable outputs, enabling robust copyright tracing in Multimodal RAG. We evaluate AQUA across diverse Multimodal RAG systems and datasets spanning different domains. The experimental results demonstrate that AQUA (1) enables the watermark images to be retrieved and reflected in the generated textual output, (2) prevents false positive retrieval from common image content, (3) remains imperceptible to users and undetectable by unauthorized filtering mechanisms, and (4) is robust to attacks such as image transformations and regeneration.
Our contribution can be summarized as follows:
@@
  • We propose AQUA, the first watermarking framework for image knowledge copyright protection in Multimodal RAG, addressing indirect watermark propagation and successful retrieval under unapparent distribution shifts and explicit watermark injection.
@@
  • We design two complementary watermarking strategies, $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$, to support more realistic black-box scenarios.
  • We conduct comprehensive experiments on two RAG datasets and RAG architectures to demonstrate the effectiveness, harmlessness, stealthiness and robustness of AQUA.
@@ 2.1 Multimodal Retrieval-Augmented Generation
Relying only on textual information is a limited approach for describing the intricate nature of the physical world. Yu et al. (2024); Mei et al. (2025); Papageorgiou et al. (2025) extend the text-only RAG framework to multimodal settings and explicitly incorporate diverse data modalities into both the retrieval and generation stages. A common strategy for enabling cross-modal retrieval is to employ powerful multimodal encoders (e.g., CLIP (Radford et al., 2021)), to map different modalities (e.g., text and images) into a shared semantic embedding space. This unification allows standard vector search algorithms like cosine similarity to retrieve relevant items across modalities based on semantic relatedness.
@@ 2.2 RAG Watermarking
Several watermarking approaches have been proposed to protect the copyright of textual knowledge in RAG. WARD (Jovanović et al., 2025) uses the LLM red-green list watermarking technology to watermark all text in the RAG knowledge base (Kirchenbauer et al., 2023; Gloaguen et al., 2024). RAG-WM (Lv et al., 2025) presents a black-box RAG watermarking approach that leverages interactions among multiple LLMs to generate high-quality watermarks. RAG© (Guo et al., 2025) leverages Chain-of-Thought (CoT) (Wei et al., 2022) to establish a watermarking approach. DMI-RAG (Liu et al., 2025) performs dataset membership inference by injecting a small number of synthetic, watermarked "canary" documents into the Intellectual Property (IP) dataset. However, existing methods on watermarking knowledge bases in RAG systems have exclusively focused on purely textual data. To the best of our knowledge, no prior work has addressed the protection of knowledge copyright in Multimodal RAG systems, particularly those integrating image and text modalities, via watermarking techniques.
@@ 3 Preliminary
In this section, we outline the workflow of the T2T Multimodal RAG system and define the notation in Section 3.1. Then, we establish the threat model for knowledge copyright protection in Multimodal RAG systems in Section 3.2.
@@ 3.1 Multimodal RAG System Workflow
The T2T Multimodal RAG system consists of three components: a retriever $\mathcal{E}$, a generator $\mathcal{G}$, and an external image knowledge base $D$. The retriever includes a text encoder $\mathcal{E}_{text}$ and an image encoder $\mathcal{E}_{img}$. Images $I_{i}$ in the external knowledge base $D=\{I_{1},\dots,I_{n}\}$ are preprocessed to a latent space through the image encoder: $e_{I_{i}}=\mathcal{E}_{img}(I_{i})\in\mathbb{R}^{d}$.
The retriever accepts the user’s text query $T$ as input, and processes it into the same latent space as images $e_{T}$$\mathcal{E}_{text}(T)\in\mathbb{R}^{d}$. Then the retriever employs a similarity function, $\text{Sim}(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\text{Score}$ (e.g., cosine similarity), to find the most relevant image knowledge according to the user’s text query: $s_{i}=\text{Sim}(e_{T},e_{I_{i}})$. Based on these similarity scores $s_{i}$, the retriever selects the top-k most relevant images as output:
@@
The original text query $T$ and the retrieved set of images $D_{retrieved}$ are combined and passed to the generator $\mathcal{G}$ to produce the final answer: $A=\mathcal{G}(D_{retrieved},T)$
@@ 3.2 Threat Model
We consider image knowledge copyright protection in Multimodal RAG services.
Defender represents the knowledge provider, aiming to detect and audit unauthorized use of its proprietary image knowledge by external Multimodal RAG services. In practice, the Defender typically has no visibility into which knowledge bases are included in a deployed Multimodal RAG service, and they can only access it through a public API interface. Defender can only operate on its own datasets to implement protection mechanisms such as injecting watermarks before contributing its data to a RaaS.
Adversary is a Multimodal RAG service provider who incorporates image datasets without authorization, with the goal of improving system performance while avoiding licensing costs. Adversary may unknowingly ingest the watermarked data and expose its presence through the system’s generated outputs, creating an opportunity for Defender to audit its misuse.
@@ 4 Methodology
AQUA is a watermarking framework designed for copyright protection of image knowledge in Multimodal RAG services, meeting four key requirements: effectiveness, harmlessness, stealthiness, and robustness. In this section, we instantiate the AQUA framework with two complementary watermarking methods, $\textbf{AQUA}_{acronym}$ and $\textbf{AQUA}_{spatial}$.
Figure 3: Illustration of the watermark injection (left) and verification (right) of AQUA.
4.1 $\textbf{AQUA}_{acronym}$
Watermark Injection. $\textbf{AQUA}_{acronym}$ addresses indirect watermark propagation from image knowledge to detectable textual output by embedding uncommon acronyms and their full names into synthetic images.
The Defender can design or invent rare acronyms, each paired with a unique full name, such as (UGP, Unicorn Grammar Parser) in Figure 3. Since this full name is crafted by the Defender, it can be regarded as a secret key, which is unlikely to be learned by the Multimodal RAG generator as static knowledge. Despite cross-modal transformation, the textual nature of the signal embedded in the image increases its chance of surviving end-to-end processing. Acronym pairs can also be generated in large quantities using an LLM (e.g., Gemini-2.5-Pro), via In-Context Learning (ICL) (Brown et al., 2020) and the prompt provided in Appendix A.1, and more examples are provided in Appendix A.2. Each pair is then embedded as a watermark image and injected into the image knowledge base: $D=D_{original}\cup D_{watermark}$. These images are designed to be minimally invasive and do not affect the model’s utility for normal queries.
Watermark Verification. In the verification phase, these acronyms are decoded through the generator’s OCR ability, producing detectable textual responses: the full name of the acronyms. Each watermark image has its own probe query $T_{probe}$ which can be used by the knowledge provider to detect unauthorized use. The $T_{probe}$ consists of two parts: a trigger $T_{trigger}$, used by the retriever to retrieve the watermark images, and an instruction $T_{instruction}$, which prompts the generator to generate the watermark-included responses that can be detected. We can formulate this construction as $T_{probe}=T_{trigger}\oplus T_{instruction}$. For example, in Figure 3, $T_{trigger}$ is “Background: UGP is a machine” and $T_{instruction}$ is “What is the full name of UGP?”. To verify the watermark signal, we define a strict substring-match protocol $\text{Eval}(\cdot,\cdot)$ based on a normalization function $\text{Norm}(\cdot)$ that lowercases and strips whitespace from both generated output $O_{RAG}$ and the verification signature $S$: